When asked about how many particles of Hydrogen are in 1 gram of Hydrogen, one student said 6.02x10^23, but another student said 5.97258x10^23. Explain the difference of opinion and say who is correct (if any).
1
Expert's answer
2019-10-29T06:44:37-0400

Molar mass of hydrogen = 2 "g\ mol^{-1}"

So particles of hydeogen in 2 gm hydrogen = "6.022\times 10^{23}" particles

So particles of hydrogen in 1 gm hydrogen = "\frac{1}{2}\times 6.022\times 10^{23}=3.011\times 10^{23}" particles

So, No one is correct .